OpenStudy (anonymous):

A man invested $22,500, some at 7% and some at 10% annual interest. The annual income from these investments was $2,130. How much did he invest at each rate?

OpenStudy (anonymous):

It's a system of equations. Can you create the system?

OpenStudy (anonymous):

The accounts are represented by \(x\) and \(y\).

OpenStudy (anonymous):

\[x+y=\dots\]

OpenStudy (anonymous):

"A man invested $22,500"\[ x+y=22500 \]

OpenStudy (anonymous):

"some at 7% and some at 10% annual interest. The annual income from these investments was $2,130." Try to create an equation with this. Hint: The equation should contain \(0.07x\) and \(0.10y\).

OpenStudy (***[isuru]***):

Hi, This is just my opinion, I think there is an easier way to solve this if u take the amount man invested at 7% interest is x then the amount he invest at 10% is (22 500 - x ) Then, as his profit is 2130 there will be a single equation like this 7x/100 + 10(22500 - x)100 = 2130 u can solve this for x and get the amount he invest at 7% in one shot and the u could use that value get other amount